OpenStudy (anonymous):

If y varies directly with x, and when y is 24, x is 6, what is y when x is 9?

OpenStudy (anonymous):

"varies directly with" lets you set up the equation y = k*x For some constant k. It's not a great equation because we don't know k yet, but it's something. "When y is 24, x is 6" lets us plug in one pair of y and x to get 23 = k*6 Which lets us find k. Once we've found k, our y =k*x equation becomes a lot better.